OpenStudy (anonymous):

f(x)=(3+9ln(x))/x i need to find the x coordinate of the absolute max...

OpenStudy (amistre64):

then lets get the derivative :)

OpenStudy (anonymous):

\[-9\ln x-6/x^2\] i think this is it

OpenStudy (anonymous):

u need help

OpenStudy (amistre64):

(x)(9/x) - (3+9ln(x))(1) ------------------- x^2 3(2 -3ln(x)) ----------- x^2

OpenStudy (anonymous):

help me

OpenStudy (amistre64):

I need help? like professional help?

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

.....me and my voices are content :)

OpenStudy (anonymous):

ok so now we need to find the absolute max by setting f' to 0

OpenStudy (amistre64):

0 will give us critical points, that we can test for minand max

OpenStudy (anonymous):

Do you have a particular bound in which you're looking for the max, or is it over (-infinity, +infinity)

OpenStudy (amistre64):

polpak is also good at this stuff :)

OpenStudy (anonymous):

x>0

OpenStudy (anonymous):

Well.. What is the limit as x approaches infinity is another thing to consider then when looking for absolute max. You have to check for critical points, but also end points. If the function goes to infinity as x goes to infinity then it has no absolute max value.

OpenStudy (amistre64):

if I see it right: 2 = 3ln(x) is what we are looking to solve

OpenStudy (anonymous):

yes that was correct...lol

OpenStudy (amistre64):

cbrt(e^2) is a possible critical point.... the only one I can come up with

OpenStudy (amistre64):

as x gets large, this goes to zero I think

OpenStudy (anonymous):

Yes, it does.

OpenStudy (amistre64):

the 2nd derivative I believe is: 3(7+6ln(x)) --------- x^3

OpenStudy (anonymous):

x=e^(3/2) would be the max value. As it approaches 0 as x goes to zero or infinity.

OpenStudy (anonymous):

i finished that problem i got a different one.... mean slope=-10 f'x=6x^2-12x-90 mean value theroem

OpenStudy (amistre64):

at cbrt(e^2) the 2nd derivative would be 33/e^2; which is positive.... I believe that it is a MIN at that point

OpenStudy (amistre64):

finished? kaputs? done? howd we do :)

OpenStudy (anonymous):

the answer was x=e^(2/3)

OpenStudy (amistre64):

hmmm.... I was right :) I like being right. even when im wrong...

OpenStudy (amistre64):

whats that new problem? cant quite decipher it

OpenStudy (anonymous):

Ok it was like another we did the average or mean slope =-10 and my f'x=6x^2-12x-90 I need to find values that make this equal to the slope

OpenStudy (amistre64):

let find the derivative that equals that slope to get the equation for the tangent line from it :)

OpenStudy (anonymous):

That is the derivative...

OpenStudy (amistre64):

ohh.... my eyes are going :) make it equal to -10; then add 10 to both sides to get it equal to 0 again and sove for x right?

OpenStudy (amistre64):

6x^2-12x-90 = -10 6x^2 -12x -80 = 0 6(x^2 -2x - 40/3) nice numbers :) remember to stop me when I do something stupid :)

OpenStudy (amistre64):

do quad formula on it to speed things up :) 144-(4)(-80)(6) 144 + 1920 12 +- sqrt(2064) --------------- right? 12

OpenStudy (amistre64):

1 +- sqrt(129)/3

OpenStudy (amistre64):

since the original will be a cubic, it stands to reason that there are 2 places that give us a slope of -10

OpenStudy (amistre64):

in fact, we can just suit up f'(x) to get that function right?

OpenStudy (amistre64):

(S) 6x^2-12x-90 dx 2x^3 -6x^2 -90x +c is in the family of curves

OpenStudy (amistre64):

we can drop the constant if we want to just use the x values to get it... right?

OpenStudy (anonymous):

\[2(3x^2 -6x -40) = 0\] \[\rightarrow 3x -6x -40 = 0\] \[\rightarrow x = \frac{6 \pm \sqrt{516}}{6}\]

OpenStudy (anonymous):

So at those x values the slope is -10.

OpenStudy (amistre64):

should be.... if I didnt mess it up :)

OpenStudy (anonymous):

Well I did it myself, so if your x matches mine it's probably right.

OpenStudy (amistre64):

I see the 1 :) can 516 get reduced?

OpenStudy (amistre64):

maybe to 129?

OpenStudy (amistre64):

close.... I mighta messed something up tho :)

OpenStudy (anonymous):

The equation is: \[f(x)=\frac{3+9 \text{Log}(x)}{x}\] The first derivative, \[f'(x)=\frac{9}{x^2}-\frac{3+9 \text{Log}(x)}{x^2} \] simplified, is: \[f'(x)=\frac{6-9 \text{Log}(x)}{x^2} \] Set the Numerator to zero and solve for x \[6-9 \text{Log}(x)=0\] \[x\text{=}e^{2/3} \] Pluggin x into f(x) gives the coordinates of the maximum funtion value: \[\left\{e^{2/3},\frac{9}{e^{2/3}}\right\}\text{ = }\{1.94773,4.62075\} \] Refer to the attachment plot of f(x).